### v1

# Mosaic IT: Enhancing Instruction Tuning with Data Mosaics

Abstract
Finetuning large language models with a variety of instruction-response pairs has enhanced their capability to understand and follow instructions. Current instruction tuning primarily relies on teacher models or human intervention to generate and refine the instructions and responses, which are costly, non-sustainable, and may lack diversity. In this paper, we introduce Mosaic Instruction Tuning (Mosaic-IT), a human/model-free method that can efficiently create rich and diverse augmentations from existing instruction tuning data to enhance the finetuned LLM.Mosaic-IT randomly concatenates multiple instruction data into one and trains the model to produce the corresponding responses with predefined higher-level meta-instructions to strengthen its multi-step instruction-following and format-following skills. Our extensive evaluations demonstrate a superior performance and training efficiency of Mosaic-IT, which achieves consistent performance improvements over various benchmarks and a $80\%$ reduction in training costs compared with original instruction tuning. Our codes and data are available at https://github.com/tianyi-lab/Mosaic-IT.
1 Introduction
The emergence of Large Language Models (LLMs) [1, 41, 39, 45, 46, 19] along with their remarkable performance in down-stream tasks [64, 55, 47, 57, 32, 34, 56, 63, 31, 4, 43, 59], has revolutionized the domains of Artificial Intelligence and Natural Language Processing. A key component of the recipe to unlock the exceptional ability of LLMs in understanding and following instructions is the technique of Instruction Tuning (IT) [38, 52, 8, 51, 62, 55], which involves the fine-tuning of LLMs on datasets comprising corresponding instruction-response pairs.
To ensure the quality of instruction tuning data, earlier efforts [1, 39, 45, 19] carefully curate extensive, diverse, and high-quality datasets manually. Although these datasets encompass a wide range of instructions to improve instruction tuning, they require the responses to be meticulously curated by human experts [20, 58, 52, 50, 14]. Alternatively, some approaches [49, 44, 54, 25, 55] leverage more capable teacher LLMs to reduce the labor-intensive process of data generation. For example, the Alpaca [44] utilizes self-instruct [49] to automatically generate diverse instruction tuning datasets, and the WizardLM [54] proposes to complicate the existing instruction data by an evolution algorithm. Building on this trend and the widely acknowledged notion that more complicated instructions are more beneficial for LLMs’ instruction-following ability [54, 65], numerous strategies [65, 53, 12, 25, 33, 24, 23, 17, 55] have been proposed to further diversify and complexify the instruction-response pairs, utilizing teacher models like ChatGPT-3.5 and GPT-4 [39].
Despite the enhanced performance in instruction-following ability offered by these existing methods, they face Two major limitations: (1) They heavily rely on teacher models to rewrite instruction-response pairs, which highlights the resource-intensive nature and their constraints on scalability; (2) They only increase the complexity within the scope of a single instruction, which limits the potential improvement in LLMs’ instruction-following capabilities. Rethinking the current methodologies, we hypothesize that the process of instruction tuning should not be constrained by one single instruction but be extended to follow several instructions at a time, which represents a higher level of instruction-following ability that is beneficial to the training process. A similar concept during the inference phase is proposed by batch prompting [5], where multiple samples are grouped in one batch allowing LLMs to generate multiple responses at one inference. However, its sub-optimal performances further indicate the complexity of this setting and the necessity of further training for this higher-level capability.
As orthogonal to the existing instruction tuning methods, we introduce Mosaic Instruction Tuning (Mosaic-IT), an innovative and model-free approach that augments existing instruction tuning datasets, which concurrently improves the LLM performances and lowers the training expenses. Specifically, in our method, multiple instructions and corresponding responses from the original dataset are concatenated into a single sample for fine-tuning, simulating the multi-instruction-following scenarios at no cost. Without applying any additional strategies, we term this simple process as the Primary Mosaic Strategy. Our empirical findings illustrate that this simple data augmentation, devoid of any supervision costs, proves remarkably effective. We posit that this mosaic strategy process significantly improves the complexity of the original instructions, learning from which directly benefits LLMs in their instruction-following ability. Additionally, this method offers the advantage of directly reducing the total count of instruction-response pairs, thereby cutting down on training iterations, and accelerating the training process significantly by approximately $80\%$ reduction.
Though effective, the Primary Mosaic strategy constrains LLMs in responding to the instructions in the original order and format, potentially limiting its further potential. Thus we further introduce three Auxiliary Mosaic Strategies aimed at enhancing the diversity and complexity of the mosaicked instruction-response pairs: Format, Permute, and Maskout, in which an additional meta-instruction is provided as a higher-level guideline for LLMs to follow the given instructions. Specifically, in the Format strategy, some arbitrary parsing formats will be defined in the meta-instruction thus forcing LLMs to follow these formats, which notably enhances the LLMs’ capacity to follow formats. In the Permutation strategy, an arbitrary permuted order is defined thus forcing LLMs to respond in a desired order. In the Maskout strategy, some arbitrary instructions are sampled which meta-instruction forces LLMs to ignore. Moreover, the use of these auxiliary strategies not only boosts the performance in several evaluation metrics but also keeps our method free of additional costs.
In summary, our primary contributions can be illustrated as follows:
  • We extend existing instruction tuning on one instruction at a time into following multiple instructions in diverse forms, which dramatically digs out the potential of the existing datasets.
  • We introduce an innovative Mosaic-IT approach for instruction tuning, supported by comprehensive evaluations showcasing its superior performance.
  • Notably, the Mosaic process is free of cost, eliminating the necessity and cost of collecting additional supervision from humans or teacher models. Moreover, our Mosaic-IT approach significantly enhances training efficiency by reducing required training iterations.
2 Related Work
Earlier research in instruction tuning primarily centered on constructing expansive, high-quality datasets through intensive curation by human experts, a process both time-consuming and labor-intensive [20, 58, 52, 50, 14]. Motivated by the success of Alpaca [44], recent studies have explored automated approaches for developing instruction-tuning datasets.
Instruction Data Improvement: WizardLM [54] first proposes an Evol Algorithm to complicate the existing data and reach supreme performance. LaMini-LM [53] innovatively generates "Topic-Guided" instructions utilizing Wiki data. Tree-Instruct [65] preliminarily explores the relationship between instruction complexity and Alignment and proposes adding nodes to complicate the instruction. UltraChat [12] establishes broad thematic scopes, systematically generating numerous instructions within each. Reflection-Tuning [25] sequentially refines both instructions and responses by focusing on specific evaluative criteria. DEITA [33] utilizes ChatGPT to diversify and then select the data. Selective Reflection-Tuning [24] proposes a teacher-student collaborative pipeline to improve and select the data. Instruction Fusion [17] proposes to utilize ChatGPT4 to merge two distinct instructions for further complexity enhancement. These advancements showcase a shift towards automating the generation and refinement of datasets, reducing reliance on human labor.
Instruction Data Selection: It is widely accepted that "quality is all you need" [46, 67] for instruction tuning. LIMA [67] demonstrates that merely 1,000 human-carefully-curated, high-quality training instances can substantially enhance the instruction-following performance. InsTag [37] employs the proprietary model, ChatGPT, to tag instruction data and select data with complex tags. Alpagasus [3] utilizes proprietary LLMs chatGPT and Claude2 to directly assess the quality of instruction tuning data. Cherry LLM [27] proposes the Instruction-Following Difficulty (IFD) scores to assess the difficulty of the instructions, which is a self-guided method in which no extra LLMs are utilized. Motivated by Humpback [28], Selective Reflection-Tuning [24] extends the IFD score to a reverse version, focusing on the feasibility of responses. [13] and [2] utilize reward models as the base scores for measuring data quality. DEITA [33] experiments on several different data selection metrics and builds a dataset with high quality. Superfiltering [26] reveals the consistency between weak and strong language models in perceiving instruction difficulty, making the filtering process much more efficient. All these works are devoted to distinguishing and selecting good data samples from bad ones for instruction tuning.
3 Methodology
Figure 1: Illustration of Mosaic-IT. The upper section represents the original Instruction Tuning method. In the middle is our method with the Primary Mosaic strategy, which concatenates instructions and corresponding responses as an augmentation for better Instruction Tuning. On the right are corresponding data examples of each method. The lower section illustrates the three Auxiliary Mosaic Strategies for Mosaic-IT.
3.1 Preliminaries
The instruction tuning dataset, defined as $D$, consists of $n$ data samples, each represented by a triplet $(\textit{Instruction},[\textit{Input}],\textit{Response})$. Historically, instruction tuning data has emphasized separate segments for both the instruction and input components to enhance control [50, 36, 44]. In contrast, contemporary datasets often integrate the input directly into the instruction for simplicity [67, 7, 54, 25, 24]. Following this trend, we define $x=\texttt{map}(\textit{Instruction},[\textit{Input}])$ as the unified instruction, and $y$ as the corresponding response. The mapping function can be as simple as concatenation, incorporating specific control tokens. Therefore, $D$ can be represented as ${(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{n},y_{n})}$, denoting a set of $n$ instruction-response pairs. Let $p_{\theta}(\cdot)$ denote the LLMs to be trained, with parameters $\theta$. In the instruction tuning setting, $p_{\theta}$ is typically fine-tuned by maximizing the following objective on each data $(x_{i},y_{i})$, $y_{i,j}$ represents the $j_{th}$ token of response $y_{i}$ and $l_{i}$ represents the token length:
$\max_{\theta}\sum_{j=1}^{l_{i}}\log p_{\theta}\left(y_{i,j}|x_{i}\right),$ (1)
3.2 Mosaic-IT
Motivated by the success of the existing data-centric instruction tuning methods, a line of approaches is proposed to further enhance the instruction-response pairs utilizing extra teacher LLMs. Though effective, all existing methods for instruction tuning restrict training samples to just one instruction, which severely limits the potential of the existing high-quality data and the instruction-following ability of the models to be trained. We hypothesize that the process of instruction tuning should not be constrained by one single instruction but be extended to follow several instructions at a time, which represents a higher level of instruction-following ability that is beneficial to the training process. Specifically, we propose Mosaic Instruction Tuning (Mosaic-IT) as shown in Figure 1, where the middle section illustrates our method with the Primary Mosaic strategy, and the lower section illustrates the three Auxiliary Mosaic Strategies.
3.2.1 Primary Mosaic strategy
Exploring the concept of concatenating unrelated instruction-response pairs into a unified instruction-response pair for training remains largely unexplored. The primary challenge lies in crafting a coherent overall instruction and acquiring the corresponding response. Most existing methods utilize a strong teacher model to rewrite and polish the instructions with prompting techniques and generate corresponding responses, introducing more cost by actually re-generating new data samples. To harness the full potential of existing data rather than directly replacing them, we introduce a simple stacking approach as shown in Figure 1 (middle), in which instructions are randomly concatenated alongside serial digits to form an overall instruction. The concatenated overall instruction is denoted as $[x_{1},...,x_{k}]$, with the corresponding overall response concatenated as $[y_{1},...,y_{k}]$. Here, $k$ denotes the number of original data samples integrated into each overall sample.
In this framework, the fundamental instruction-following capability is grounded in the existing instruction-response pairs, and the mosaic strategy extends this capability to a higher level in which LLMs are forced to follow multiple instructions. It represents a much more complicated scenario that benefits LLMs compared with traditional single-task instructions. Consequently, the objective function for each concatenated overall data sample can be formulated as follows:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y_{1},...,y_{k}]_{j}|[x_{1},...,x_{k}]\right),$ (2)
Here, $[y_{1},...,y_{k}]_{j}$ denotes the $j{th}$ token of the overall response, and $l$ represents its length. This formulation encapsulates the essence of our approach, optimizing the model parameters $\theta$ to maximize the likelihood of generating the correct sequence of responses for the given overall instruction.
3.2.2 Auxiliary Mosaic Strategies
Though effective, this simple Primary Mosaic strategy constrains LLMs in responding to the instructions with the original order and format, potentially limiting its generalization and practical usage. In our method, the instructions and corresponding responses from the original dataset can be viewed as atomic components and our method randomly combines these elements together to form new instructions and responses. This nature allows us to further complicate this process with fancier strategies thus forcing LLMs to follow more complicated overall instructions. Hence, we propose three Auxiliary Mosaic Strategies to complicate and diversify the mosaicked samples as shown in Figure 1 (lower), including Format, Permute, and Maskout, with meta-instructions guiding them.
Format In the Format strategy, some arbitrary formats are defined in the meta-instruction to force LLMs to follow these formats. The formats mainly contain two categories: 1) Serial Digit Format and (2) Response Parsing Format. The serial digits establish the initial instruction order that guides LLMs to follow sequentially. We manually define $10$ types of serial digit format, which will be randomly sampled during each mosaic process. For response parsing, we simulate the scenario where the users try to extract specific information from the responses. We define $27$ types of parsing brackets and $17$ types of parsing text pairs, which will be randomly sampled and assembled during each mosaic process. Examples can be found in Appendix C, which can be easily extended for customized training settings. We denote responses with specific formats as $y^{\prime}_{i}=wrap(y_{i},s_{format})$, and $l$ as the token length of the overall response. An additional meta-instruction $s_{format}$ specifying the required format will be included in the overall instruction. Thus, the objective function for each mosaic data point:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y^{\prime}_{1},...,y^{\prime}_{k}]_{j}|[x_{1},...,x_{k},s_{format}]\right),$ (3)
Permute and Maskout Building upon the Format strategy, we further introduce two strategies for our Mosaic-IT, Permutation and Maskout.
In the Permute strategy, an arbitrary permuted order is defined in the meta-instructions, forcing LLMs to follow. Moreover, several high-level rules are defined to ensure the complexity and diversity of meta-instructions, e.g., forcing LLMs to respond to each instruction in the randomly generated permutation list, forcing LLMs to respond in the alphabetical order of each instruction, forcing LLMs to respond according to the length of instructions, etc. The detailed rule types and descriptions are depicted in Appendix C. These various meta-instructions not only provide higher-level guidelines for LLMs to follow multiple instructions but also inherently enhance the instruction perception ability of LLMs. The meta-instruction is denoted as $s_{permute}$ and is included in the overall instruction. The permuted response list is denoted as $[y^{\prime}_{1^{\prime}},...,y^{\prime}_{k^{\prime}}]=Permute([y^{\prime}_{1},...,y^{\prime}_{k}],s_{permute})$. Thus the objective function can be formulated as below:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y^{\prime}_{1^{\prime}},...,y^{\prime}_{k^{\prime}}]_{j}|[x_{1},...,x_{k},s_{format},s_{permute}]\right),$ (4)
In the Maskout strategy, some arbitrary instructions are selected in the meta-instructions forcing LLMs to ignore them. Several high-level rules are also defined similarly to the permute strategy, including forcing LLMs to ignore the instructions with given random digits, forcing LLMs to ignore the longest one/several instructions, forcing LLMs to ignore odd-numbered instructions, etc. The details are provided in Appendix C. Similarly, the meta-instruction is denoted as $s_{maskout}$ and the response list is denoted as $[y^{\prime}_{1},...,y^{\prime}_{m}]=Maskout([y^{\prime}_{1},...,y^{\prime}_{k}],s_{maskout})$, where $m$ is the count of responses after masking out. Thus the objective function can be formulated as below:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y^{\prime}_{1},...,y^{\prime}_{m}]_{j}|[x_{1},...,x_{k},s_{format},s_{maskout}]\right)$ (5)
It’s important to note that our mosaic strategies entail no supervision cost, and the predefined rules are flexible and have the potential for further extension. We utilize the version with three auxiliary strategies as our default Mosaic-IT.
How to decide the Number of Instructions $k$: Number of Instructions denotes the number of original data samples that are integrated into an overall sample. In addition to the detailed mosaic strategies being used, this count also dramatically affects the effect of Mosaic-IT. Our experiments reveal that larger and more diverse numbers of instructions will benefit LLM training. By default, we set the maximum number of instructions as $k_{max}=10$, and randomly sample an integer that is smaller or equal to $k_{max}$ under a uniform distribution. If the number causes the data sample to be longer than the max length, it will be automatically reduced to the max number which remains the sample length within the limits.
4 Experimental Setup
4.1 Implementation Details
The experiments are conducted on three base pre-trained models: Llama2-7B, Llama2-13B [46], and Mistral-7B [19]. The training datasets and detailed training configurations are introduced in Appendix B.
4.2 Evaluation Metrics
Pair-wise Comparison by using powerful LLMs like GPT-4 is recently widely accepted and becoming a common practice [46, 7, 11, 35, 6]. The evaluation of responses from LLMs, especially in open-domain contexts where definitive ground truth is hard to establish, continues to be an intricate and evolving research domain. Recent studies, however, have indicated a notable alignment between GPT-4’s performance evaluations and human assessments [66, 29, 42], thereby establishing a credible foundation for this evaluative methodology. We adopted test instruction sets from WizardLM [54], comprising 218 diverse, human-curated instructions for pair-wise comparison. We directly follow the evaluation framework proposed by [3, 27], which evaluates responses on a scale spanning from 1 to 10 across multiple dimensions. To further address positional bias, as discussed by [22, 48], the comparison is conducted in two distinct sequences, LLM1’s response first and then LLM2’s response first, ensuring a fair assessment of model performance. Evaluation outcomes are categorized into ’win-tie-loss’ for each instruction. The detailed evaluation prompt is provided in Appendix D.
Human Evaluation is further implemented to substantiate the superiority of our approach based on the WizardLM test set. The test set contains $100$ samples randomly sampled from the original WizardLM test set. Three human evaluators were tasked with comparing the outputs generated by the models under consideration, using the same criteria as in the previous pairwise evaluation. Each evaluator was presented with three response options: Win, Tie, and Loss. The final outcomes were determined by a majority vote. More evaluation metrics including Open LLM Leaderboard, Alpaca-Eval Leaderboard, MT-Bench and IF Eval are introduced in Appendix B.
5 Experimental Results
5.1 Main Results
In this section, we present the evaluation results comparing our methods with the baseline methods on Three baseline models (Mistral-7B [19], Llama2-7B [46], Llama2-13B) and Three instruction tuning datasets (Alpaca-GPT4 [40], Alpaca [44], WizardLM-70k [54]), on Two general evaluation settings (Pair-Wise Comparison and Open LLM leaderboard) described above, as shown in the Table 1. Pair-Wise Winning Score indicates the result directly comparing our models with the corresponding baseline models, which is calculated as (Num(Win)--Num(Lose))///Num(All) $+1$. These values that are greater than $1.0$ represent better responses generated by our models. The performances on the Huggingface Open LLM Leaderboard are also presented, and we bold the greater average values for each comparison. The consistent outperforming results on different base models and datasets represent the effectiveness and robustness of our methods. Moreover, comparing the results on different base models, the performance growths of our method on Mistral-7B and Llama2-13B models are greater than the Llama2-7B models, indicating that our method works better on relatively better base models. We hypothesize it is because our method largely increases the complexity of instructions by proposing a higher level of meta-instructions while keeping the quality of original responses unchanged, making it hard for weaker models to learn.
To better understand how our method improves the instruction-following abilities of LLMs, we further compare the performance on other Three benchmarks for fine-grained analysis based on the Mistral-7B base model with two datasets as shown in Table 2. On the MT-Bench, the 1-round scores of our method are higher while the 2-round scores are slightly lower, indicating that our method mainly improves the response quality for single-round conversations, which is reasonable as the meta instructions only focus on single-round formats. On the Alpaca Eval benchmark, our method has a consistent improvement with or without the Length Control (LC), indicating that the improvement of response qualities does not directly originate from the length of responses. On the IF Eval benchmark, our method consistently improves the performances on all 4 different settings, both Prompt-level and Instruction-level, both Strict version and Loose version. Compared with the previous benchmarks, IF Eval mainly focuses on the format-following ability of LLMs. The consistent improvement in this benchmark represents that our method not only improves the response qualities of the LLMs but also improves their controllability regarding formats. Given that our method is a cost-free augmentation technique that does not rely on any additional models, the observed improvements are remarkable.
Further Human Evaluations are conducted on Mistral-7B with Alpaca-GPT4 and WizardLM dataset. For the comparison on (1) Alpaca-GPT4: the model using Mosaic-IT wins on $68$ out of $100$ instruction, ties on $3$, and losses on $29$ instructions; on (2) WizardLM: the model using Mosaic-IT wins on $63$ out of $100$ instruction, ties on $6$, and losses on $31$ instructions. This human evaluation also further verifies the effectiveness of our Mosaic-IT.
Model Dataset Method Pair-wise $\uparrow$ Huggingface Open LLM Leaderboard $\uparrow$
Winning Score Average ARC HellaSwag MMLU TruthfulQA
Mistral-7B Alpaca-GPT4 Baseline 1.000 59.70 55.03 78.87 56.01 48.88
Mosaic-IT 1.349 63.65 59.04 81.85 60.09 53.62
Alpaca Baseline 1.000 55.15 51.96 74.61 52.85 41.20
Mosaic-IT 1.390 58.86 56.23 79.57 57.06 42.58
Wizard-70k Baseline 1.000 57.86 51.88 77.93 53.76 47.89
Mosaic-IT 1.161 61.11 57.85 82.13 57.42 47.08
Llama2-7B Alpaca-GPT4 Baseline 1.000 58.71 54.69 80.05 47.89 52.21
Mosaic-IT 1.073 58.84 54.18 80.54 47.92 52.70
Alpaca Baseline 1.000 55.25 54.35 78.65 47.02 40.98
Mosaic-IT 1.096 55.32 53.75 78.65 46.88 41.98
Wizard-70k Baseline 1.000 57.09 54.18 79.25 46.93 48.02
Mosaic-IT 1.197 57.41 54.69 79.69 48.11 47.13
Llama2-13B Alpaca-GPT4 Baseline 1.000 61.47 58.70 83.12 54.13 49.92
Mosaic-IT 1.110 63.26 58.87 83.54 55.75 54.87
Alpaca Baseline 1.000 57.63 57.25 81.23 54.13 37.91
Mosaic-IT 1.046 58.80 56.57 81.79 54.28 52.55
Wizard-70k Baseline 1.000 61.24 57.04 83.39 55.76 48.78
Mosaic-IT 1.078 61.50 58.70 83.69 56.44 47.18
Table 1: The performance comparison on the Pair-wise Comparison Winning Score and the Open LLM Leaderboard, on 3 different base models and 3 different instruction tuning datasets.
Dataset Method MT-Bench $\uparrow$ Alpaca Eval 2 $\uparrow$ IF Eval $\uparrow$
1-round 2-round Rate Rate (LC) Prompt (S) Inst (S) Prompt (L) Inst (L)
Alpaca-GPT4 Baseline 6.44 5.26 3.98 7.28 32.53 42.93 35.86 45.92
Mosaic-IT 7.11 4.69 5.00 7.81 37.15 48.56 38.08 50.23
Wizard-70k Baseline 6.21 4.70 4.13 6.46 39.56 49.88 41.96 53.00
Mosaic-IT 6.95 4.32 4.44 7.56 40.85 51.80 45.47 56.47
Table 2: The performance comparison on the MT-Bench, Alpaca Eval, and IF Eval Benchmarks. Rate(LC) in Alpaca Eval represents the length-controlled win rates. In IF Eval, Prompt and Inst represent Prompt-level and Instruction-level accuracy; S and L represent Strict and Loose versions.
5.2 Ablation studies
In this section, extensive experiments are conducted on Mistral-7B using with the Alpaca-GPT4 dataset to verify the effectiveness of our method. We utilize Pair-wise comparison for evaluation.
Winning Score Win Tie Lose
Primary 1.261 110 55 53
Format 1.284 109 62 47
Permute 1.334 118 55 45
Maskout 1.376 121 58 39
Permute/Maskout 1.349 123 48 47
(a) Ablation on Mosaic-IT strategies.
Winning Score Win Tie Lose
Max Count = 2 0.989 70 75 73
Max Count = 4 1.142 92 65 61
Max Count = 6 1.303 111 62 45
Max Count = 8 1.294 112 58 48
Max Count = 10 1.349 123 48 47
Max Count = 12 1.376 124 52 42
(b) Ablation on the Max Number of Instructions.
Table 3: Ablation on (a) Mosaic-IT strategies and (b) Max Number of Instructions.
Ablation on Mosaic Strategies is presented in Table 3(a). Primary represents the Primary Mosaic Strategy. The winning score of this setting is greater than $1.0$, indicating a better performance compared with the baseline method. This comparison directly verifies the effectiveness of the idea of introducing multiple instructions during training, which complicates the instructions at no cost and improves the instruction-following ability of LLMs. Format represents the Format Strategy. Although the winning score is only slightly greater than the naive version, this version makes it possible for LLMs to follow the customized user-defined formats, indicating great potential for the controllability of LLMs. Moreover, the format version can be easily used with other types of meta instructions, showing great extensibility. Permute represents the Permute Strategy that builds on the Format Strategy with a probability of $\frac{1}{2}$, similar to Maskout. Permute/Maskout represents our default setting, where the Permute or Maskout Strategies are utilized together with the Format Strategie with a probability of $\frac{1}{3}$. All these $3$ settings show higher performance than the format version, indicating the effectiveness of Auxiliary Mosaic Strategies which define more complicated meta instructions.
Ablation on the Max Number of Instructions is presented in Table 3(b), including the pair-wise comparison values. As shown in the table, when the max number is set as $2$, i.e. at most $2$ instructions/responses are concatenated together, the performance is almost the same as the baseline, indicating the ineffectiveness. However, when the max number grows, the corresponding winning scores also grow consistently. This trend shows that the more instructions concatenated together, the better the instruction-following ability. We hypothesize that, with the growth of the number of instructions, the overall instruction becomes much harder to follow, especially for the permute and maskout strategies, which benefits LLMs’ instruction-following capability.
Winning Score Win Tie Lose Mix $\leq$ 5
Fix 0.982 90 34 94 2.39%
Exponential 0.995 94 29 95 2.58%
Pareto 1.417 129 51 38 8.94%
Log-normal 1.431 136 40 42 6.83%
Logistic 1.417 123 49 46 15.84%
Uniform 1.349 123 48 47 51.45%
Table 4: Ablation on the Distribution of Number of Instructions, including the pair-wise comparison values (left) and the visualization of distribution comparisons (right). The detailed numbers of instructions for different settings can be shown in Appendix A. “Mix $\leq 5$” represents the percentage of samples with the number of instructions less or equal to $5$.
Ablation on the Distribution of Number of Instructions is presented in Table 4, including the pair-wise comparison values (left) and detailed number distribution comparisons (right), which aims at identifying how this count distribution affects the performance of our method. The detailed distribution formula, and data counts are provided in the Appendix A. Fix represents the setting where all the overall instructions are concatenated with a fixed number of instructions, which we set as $10$ unless the overall instructions exceed the max length limit. Exponential represents the setting where the number of instructions is sampled following the exponential distribution. Under these two settings, less than $3\%$ of the overall instructions are concatenated by less or equal to $5$ original instructions. The lack of few-instruction concatenated samples negatively affects the LLMs’ ability to follow the single instruction, which is employed by most of the existing evaluation methods, leading to worse performances. Pareto, Log-normal, and Logistic represents the corresponding distribution that are utilized for sampling. Different from the above two settings, approximately $10\%$ of the overall instructions are composed of fewer original instructions, thus ensuring the LLMs are trained with samples with sufficiently diverse lengths, resulting in optimal performances. Uniform is our default setting, representing using the uniform distribution where different numbers are sampled evenly. In this situation, the LLMs are trained with samples with the most diverse lengths, thus avoiding the LLMs overfit to simple lengthy responses.
6 Further Discussion
Figure 2: The training loss curve comparisons between the original instruction tuning process and our Mosaic-IT with w datasets on (a) Mistral-7B, (b) Llama2-7B, and (c) Llama2-13B. Two conclusions can be drawn from the comparisons: (1) Much fewer training iterations are required in our method, indicating the more efficient training process; (2) The “stair-like” loss curves for the original training process indicate a potential memorizing effects, while our loss curves are more smooth. All the training settings are kept the same between the baseline models and Mosaic-IT models, including the Learning Rate, Warm-up Ratio, Learning Rate Schedule (Cosine), Batch Size, etc.
6.1 Memorizing
In the original instruction tuning process, each data sample will be trained several times for LLMs without many changes to the instructions and responses. Though effective, this training process poses risks to the potential memorizing effects on training samples, which can be partially indicated by the “stair-like” training loss curves as shown in Figure 2. In the figure, all the training settings are kept the same between the baseline models and Mosaic-IT models, including the Learning Rate, Warm-up Ratio, Learning Rate Schedule (Cosine), Batch Size, etc. For the baseline methods, the training loss hardly decreases within each epoch of training but drops dramatically when the LLMs meet the same training samples again, which indicates a potential memorizing effect of training samples and potential overfitting. However, when utilizing our method, the random mosaics of original instructions with diverse and complex meta-instructions largely diversify the overall training instructions. Although each original data sample will still be seen by LLMs several times during training, the overall context varies dramatically as each original sample is only an atomic element of the overall mosaic sample, indicating that there will be no identical overall instructions during the whole training process. Thus this augmentation largely alleviates the potential memorizing and overfitting problems as shown in the figure, where the training loss decreases smoothly, representing the gradual learning process.
6.2 Efficiency
Settings Baseline Fix Exponential Pareto Log-normal Logistic Uni-2 Uni-4 Uni-6 Uni-8 Uni-10 Uni-12
Time (min) 827 121 129 133 133 143 716 426 305 245 202 173
Time Ratio 100.0% 14.6% 15.6% 16.1% 16.1% 17.3% 86.6% 51.5% 36.9% 29.6% 24.4% 20.9%
Winning Score 1.000 0.982 0.995 1.417 1.431 1.417 0.989 1.142 1.303 1.294 1.349 1.376
Table 5: The training time comparison of different settings, and the pair-wise winning scores are also provided for better illustration. “Uni-2” represents uniform distribution with max count as $2$. Mosaic-IT can reduce the training time to approximately $16\%$ to $25\%$ while achieving better performance.
One of the benefits of our method is the efficiency of the training process. Given an existing dataset, our mosaic processes largely decrease the number of total overall instructions and the total number of gradient descents, leading to a reduction in the training process. The detailed comparison is shown in Table 5, which is based on the Mixtral-7B model on the Alpaca-GPT4 dataset. The time is calculated based on four NVIDIA A100 Graphic Cards. As shown, our method greatly decreases the training time to approximately $16\%$ to $25\%$ while achieving better performances, especially when there are mosaic samples with larger permutation counts.
6.3 Limitation
The potential limitations of our work: (1) Currently, three Auxiliary Mosaic Strategies with corresponding high-level rules are proposed and utilized in our method, however, we believe more strategies and predefined rules can be further introduced. (2) The optimal distribution of the number of instructions for the mosaic process still needs further justification in future studies. (3) It is unknown whether the inclusion of extra models or careful curation/selection of instructions for concatenation will further improve the performance of Mosaic-IT largely.
7 Conclusion
We introduce Mosaic Instruction Tuning (Mosaic-IT), a novel, human/model-free method to enhance instruction tuning for LLMs. By concatenating multiple instruction-response samples and using higher-level meta-instructions, Mosaic-IT improves multi-step and format-following capabilities. Our evaluations show superior performance and an $80\%$ reduction in training costs compared to the original methods. Mosaic-IT’s simplicity and efficiency make it a scalable solution for improving LLMs without extensive human intervention or resource-intensive teacher models. Our results highlight the potential of innovative data augmentation techniques in advancing LLM capabilities.
Appendix A Detailed Distribution for Ablation on Mixture Distribution
A.1 Distribution description
The detailed distribution descriptions and formulas are provided below.
Exponential Distribution [^1] https://numpy.org/doc/stable/reference/random/generated/numpy.random.exponential.html: The exponential distribution is a continuous probability distribution used to model the time or space between events in a Poisson process. The probability density function (PDF) of the exponential distribution is:
$f(x;\lambda)=\lambda e^{-\lambda x}\quad\text{for }x\geq 0,$
where $\lambda=1$ by default in our setting. We will resample with this distribution if the sampled value $x_{sample}$ is greater then $k_{max}$.
Log-normal Distribution [^2] https://numpy.org/doc/stable/reference/random/generated/numpy.random.lognormal.html: The log-normal distribution is a continuous probability distribution of a random variable whose logarithm is normally distributed. It is often used to model variables that are positively skewed, such as income, stock prices, and other financial data. The probability density function (PDF) for a log-normal distribution is given by:
$f(x;\mu,\sigma)=\frac{1}{x\sigma\sqrt{2\pi}}\exp\left(-\frac{(\ln x-\mu)^{2}}{2\sigma^{2}}\right)\quad\text{for}\quad x>0$
where $\mu=0$ and $\sigma=0$ by default in our setting. We will resample with this distribution if the sampled value $x_{sample}$ is greater then $k_{max}$.
Logistic Distribution [^3] https://numpy.org/doc/stable/reference/random/generated/numpy.random.logistic.html: The logistic distribution is a continuous probability distribution used in various fields, including logistic regression, modeling growth, and in some cases as an alternative to the normal distribution due to its heavier tails. The probability density function (PDF) for the logistic distribution is given by:
$f(x;\mu,s)=\frac{e^{-(x-\mu)/s}}{s\left(1+e^{-(x-\mu)/s}\right)^{2}}$
where $\mu=0$ and $s=2$ by default in our setting. We will resample with this distribution if the sampled value $x_{sample}$ is greater then $k_{max}$.
Pareto Distribution [^4] https://numpy.org/doc/stable/reference/random/generated/numpy.random.pareto.html: The Pareto II or Lomax distribution is a shifted Pareto distribution. It can be considered as a simplified version of the Generalized Pareto distribution, with the scale set to one and the location set to zero. The probability density function (PDF) for the Pareto distribution is:
$f(x;\alpha)=\frac{\alpha m^{\alpha}}{x^{\alpha+1}}\quad\text{for}\quad x\geq m,$
where $m=1$ and $\alpha=1$ by default in our setting. We will resample with this distribution if the sampled value $x_{sample}-1$ is greater then $k_{max}$.
After getting $x_{sample}$, a floor function will be utilized to get the corresponding integer and the final concatenation count $k=k_{max}-floor(x_{sample})$.
A.2 Distribution visualization
The detailed data counts for different distribution are provided in Figure 3.
(a) Fix Max Number
(b) Exponential Distribution
(c) Log-normal Distribution
(d) Logistic Distribution
(e) Pareto Distribution
(f) Uniform Distribution
Figure 3: Bar plots of detailed data counts for different distributions in the Ablation on the Numbers of Instructions: (a) Fix Max Number, (b) Exponential Distribution, (c) Log-normal Distribution, (d) Logistic Distribution, (e) Pareto Distribution, (f) Uniform Distribution.
Appendix B Experimental Setup
B.1 Implementation Details
The experiments are conducted on three base pre-trained models: Llama2-7B, Llama2-13B [46], and Mistral-7B [19]. We utilize the prompt and code base from Vicuna [7] and flash attention [10] while the overall training arguments are aligned with the common training configuration. The Adam optimizer [21] is utilized with the batch size of $128$ and with the max token length of $2048$. When training the baseline models on Llama2-7B and Llama2-13B, the maximum learning rate is set to $2\times 10^{-5}$ with the warmup rate as $0.03$ for $3$ epochs. When training the baseline models on Mistral-7B, the maximum learning rate is set to $1\times 10^{-5}$ with the warmup rate as $0.1$ for $3$ epochs. When training with Mosaic-IT, we run the mosaic process $4$ times for each experiment to simulate $4$ epochs of training. Then these augmented data are mixed together used for training $1$ epoch while all other configurations are kept the same as baselines.
B.2 Training Dataset
The Alpaca dataset [44] comprises $52,000$ instruction-following samples and is constructed utilizing the self-instruct paradigm [49]. This dataset was produced by employing OpenAI’s text-davinci-003 model. Characterized as a classical dataset with moderate quality attributes, the Alpaca dataset serves as an initial platform to validate our methodology. To further substantiate our approach using a dataset of inherently high quality, we also applied our method to the Alpaca-GPT4 dataset [40], which features responses generated by GPT4. The WizardLM dataset [54] is also utilized in our method, which contains $70,000$ samples created by the evolution algorithm proposed by them. With ChatGPT-3.5 utilized, the data quality on WizardLM is largely guaranteed.
B.3 Evaluation Metrics
Open LLM Leaderboard, employing the evaluation framework from Eval Harness [16], offers a detailed and systematic approach to assessing the capabilities of generative language models through a set of diverse evaluation tasks. This methodology zeroes in on four pivotal benchmarks: ARC [9], HellaSwag [60], MMLU [18], and TruthfulQA [30]. These benchmarks collectively provide a comprehensive evaluation of the models’ reasoning abilities, their grasp of common-sense knowledge, and their accuracy in presenting factual information. Consequently, the leaderboard presents valuable insights.
Alpaca-Eval Leaderboard, leveraging the AlpacaFarm evaluation dataset, presents a dependable and efficient automated evaluation tool for LLMs [29, 15]. This tool benchmarks the responses generated by LLMs against those from Davinci003, focusing on the models’ ability to adhere to generic user instructions.
MT-Bench (Multi-turn Benchmark) [66] is a benchmark tool designed for automated evaluating LLMs in multi-turn dialogue settings. It focuses on analyzing conversation flow and the model’s ability to follow instructions with 80 high-quality, multi-turn questions.
IFEval (Instruction-Following Eval) [61] is a straightforward and easy-to-produce evaluation benchmark focusing on a set of “verifiable instructions”. It contains 25 types of verifiable instructions and 541 prompts, with each prompt containing one or multiple verifiable instructions.
Appendix C Predefined Rules
Examples of predefined formats can be found in Table 6 and detailed predefined rule descriptions can be found in Table 7.
Serial Digit Parsing Bracket Parsing Text Assembled Examples
i (text) BEGIN, END 1. (BEGIN)response(END)
(i) [text] START, END (1). [START]response[END]
[i] <text> RESPONSE, END [1]. <RESPONSE>response<END>
<i> <<text>> RESPONSE, END OF RESPONSE <1>. <<RESPONSE>>response<<END OF RESPONSE>>
<<i>> |text| OPEN, CLOSE <<1>>. |OPEN|response|CLOSE|
###i [|text|] OPEN RESPONSE, CLOSE ###1. [|OPEN RESPONSE|]response[|CLOSE|]
##i <|text|> INITIATE, TERMINATE ##1. <|INITIATE|>response<|TERMINATE|>
##i## #text# START POINT, END POINT ##1##. #START POINT#response#END POINT#
|i| *text* RES_START, RES_END |1|. *RES_START*response*RES_END*
||i|| @text@ RES, /RES ||1||. @RES@response@/RES@
Table 6: Examples of predefined formats, including the Serial Digit formats and Response Parsing formats. “i” represents the real number serial number, “text” represents the replaceable parsing text, and “response” represents the real response of the concatenated overall instructions/responses. The response parsing formats are composed of the parsing bracket and text. In each mosaic process, random formats will be sampled simulating the real-world user-defined formats. The last column represents the assembled examples using the formats in the same row.
Strategy Rule Name Rule Description
Permute FIX Respond in the order of a provided list.
Permute REVERSE Respond in reverse of the original order.
Permute ALPHA Respond in the alphabetical order of the first letter of tasks.
Permute REVERSE_ALPHA Respond in the reverse alphabetical order of the first letter of tasks.
Permute LENGTH_WORD Respond according to the length (words) of tasks, respond to short ones first.
Permute REVERSE_LENGTH_WORD Respond according to the length (words) of tasks, respond to long ones first.
Permute LENGTH_CHAR Respond according to the length (characters) of tasks, respond to short ones first.
Permute REVERSE_CHAR_WORD Respond according to the length (characters) of tasks, respond to long ones first.
Permute ODD_EVEN First respond to the odd-numbered tasks, then the even-numbered ones.
Permute EVEN_ODD First respond to the even-numbered tasks, then the odd-numbered ones.
Maskout FIX Ignore the tasks provided in the list.
Maskout WORD_LONG Ignore the longest one/several task(s) according to the word count.
Maskout WORD_SHORT Ignore the shortest one/several task(s) according to the word count.
Maskout ODD Ignore the odd-numbered tasks.
Maskout EVEN Ignore the even-numbered tasks.
Table 7: Predefined rules for the Permute and Maskout strategy. A random rule will be sampled for each mosaic process, which largely complicates and diversifies the mosaicked instructions.
Appendix D Prompt for Evaluation
The detailed pair-wise comparison prompt for the pair-wise comparison is in Figure 4.
  Prompt for Performance Evaluation
  System Prompt
You are a helpful and precise assistant for checking the quality of the answer.
User Prompt
[Question]
Question
[The Start of Assistant 2’s Answer]
Answer 2
[The End of Assistant 2’s Answer]
[The Start of Assistant 2’s Answer]
Answer 2
[The End of Assistant 2’s Answer]
We would like to request your feedback on the performance of two AI assistants in response to the user question displayed above.
Please rate the helpfulness, relevance, accuracy, level of details of their responses. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance.
Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment.
Figure 4: The prompt we used to request GPT4-Turbo to evaluate the responses.